OpenStudy (anonymous):

please helpppp! I need help so bad :( 8. What is the slope of the line that passes through the pair of points (-7. 1, 11.4) and (-3. 1, 3.4) ? a. -1 b. -2 c. 1 d. 2

OpenStudy (anonymous):

@BunnyBree :3 help mee

OpenStudy (anonymous):

Alright, so here's the equation. m = y2 - y1/x2 - x1 I'll tell you where to place the numbers, but calculate it by yourself. Let me know what you get, and I'll tell you if you got it right. (: For y2 and y1: 3.4 - 11.4 For x2 and x1: -3.1 - (-7.1) Calculate these and then divide them. So, 3.4 - 11.4/ -3.1 - (-7.1)

OpenStudy (anonymous):

okiii

OpenStudy (anonymous):

let me just try calculating it ^_^

OpenStudy (anonymous):

wait, was I suppose to put calculate all that? cause I got 6.822581

OpenStudy (anonymous):

Okay, I'll simplify it. 3.4 - 11.4 = -8 -3.1 - (-7.1) = 4 Divide -8 by 4. What do you get?

OpenStudy (anonymous):

phhh!

OpenStudy (anonymous):

ohh*

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

I got -2 ^_^